# TableMaster: A Recipe to Advance Table Understanding with Language Models

### v1

@@ 5.2 Main Results
As shown in Table 1, our TableMaster approach consistently achieves the highest performance across both WikiTQ and TabFact under different backbone models (gpt-3.5-turbo, gpt-4o-mini, and Llama-3.1-70B). On WikiTQ, TableMaster outperforms the strongest baselines by +3.51, +13.40, and +12.39 points, respectively. A similar trend is observed on TabFact, with improvements of +1.73, +1.19, and +4.10 points, demonstrating the robustness of our method across diverse large language models. Results on the FetaQA dataset are provided in Appendix C. These results confirm that TableMaster not only generalizes well across different base language models but also significantly enhances table understanding and reasoning in complex QA tasks.
@@ 5.3 Ablation Study
These results reveal that TableMaster ’s performance relies on a combination of structural understanding, content adaptation, and an effective reasoning strategy. Among these, structural extraction and reasoning components contribute the most to performance, with textual reasoning being the most critical. Additionally, while content understanding components such as re-construction and verbalization provide meaningful improvements, their impact varies across datasets. These findings validate the effectiveness of our framework in handling diverse table understanding tasks.
@@ 6 Conclusion
In this paper, we explore table understanding with language models. Given the characteristics of tabular data, we identify key challenges in table understanding. To overcome these challenges, we propose TableMaster, a recipe and comprehensive framework that integrates multiple solutions. Extensive analyses and experiments demonstrate our findings and the effectiveness of TableMaster. In the future, we plan to extend and refine the framework to improve its performance across diverse practical applications.
Acknowledgments
This work was completed during my research internship. I would like to express my gratitude to my mentor and the friends I made during this time. I am also grateful to the company for providing the computing and API resources that made this work possible.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Limitations, Extendability, and Future Works
@@
A.1 Technical Refinement
Wild Table. In our experiments, the tables in the three datasets we use are already cleaned; therefore, we do not explicitly implement table normalization in our evaluation experiments. However, we conduct analysis experiments to highlight the importance of table normalization for handling wild tables. In practical scenarios, various tools are available for table normalization. Regular expression matching can be employed for formatting, and small language models can also be leveraged to efficiently process and normalize tables (Nahid & Rafiei, 2024b).
@@ A.1 Technical Refinement
Table-of-Focus Construction. In designing the Table-of-Focus, we employ two efficient methods: LM prompting for column lookup and SQL generation for row lookup. The Table-of-Focus is then constructed based on the results of these two lookups. Many previous works (Ji et al., 2024; Wang et al., 2024) have introduced complex approaches for extracting relevant sub-tables. In contrast, our method remains intentionally simple, prioritizing efficiency and adaptability. We believe that in the future, more advanced techniques may emerge to further enhance the extraction of key information.
Table Verbalization. To facilitate the implementation of TableMaster, we utilize language models themselves to verbalize the table. However, the quality of the generated text is not optimal due to the challenges of open-ended text generation. Several existing studies, such as Table-to-Text (Parikh et al., 2020), have explored this sub-task. In the future, we can enhance performance and efficiency by replacing this step with specifically trained small language models, which could further improve the semantic density of the verbalized table.
Adaptive Reasoning. Adaptive reasoning can be unstable, as language models may not always select the optimal strategy. We further explore this issue in Appendix J. In the future, training a dedicated machine learning model to guide LMs in selecting the most effective reasoning strategy could improve stability and performance.
Information Missing. The construction of the Table-of-Focus involves a trade-off between precision and recall. If recall is insufficient, essential information may be missing for final reasoning, while low precision can render the extracted content less useful. Although we use re-construction to mitigate information loss during the Table-of-Focus construction process, our analysis reveals that some information missing persist in row lookup. We further investigate this issue in Appendix K.
Efficiency. Efficiency is crucial in table processing and table understanding. To enhance efficiency, we incorporate the table peek technique, which reduces the context that language models need to process at certain steps. We further explore this technique in Appendix F and analyze the overall efficiency in Appendix G. In real-world applications, for optimal efficiency, we consider replacing certain steps with specialized small language models, balancing the trade-off between efficiency and performance .
A.2 Downstream Applications
@@
Appendix B Detailed Settings of Challenge Analysis Experiments
We conduct extensive experiments to analyze the challenges of table understanding with language models (LMs). Specifically, we perform challenge analysis experiments on the WikiTQ dataset (Pasupat & Liang, 2015), which consists of 4,344 data instances. Following previous work (Wang et al., 2024; Liu et al., 2024b), we use the exact match of the final answer as the evaluation metric to measure accuracy. Our experiments utilize OpenAI models hosted on Microsoft Azure. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. For each model, we use the following versions: gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), and o1 (o1-preview-0912).
@@ Appendix B Detailed Settings of Challenge Analysis Experiments
Appendix C Experimental Results on the FetaQA Dataset
@@
Fine-Tuning (T5-large) (Ye et al., 2023) 30.54 63 41 53
End-to-End QA (Codex) (Chen et al., 2021) 27.96 62 40 52
End-to-End QA (PaLM 2) (Wang et al., 2024) 28.37 63 41 53
Dater (PaLM 2) (Ye et al., 2023) 29.47 63 41 53
Chain-of-Table (PaLM 2) (Wang et al., 2024) 32.61 (+14.9%) 66 (+4.8%) 44 (+7.3%) 56 (+5.7%)
@@
PaLM 2 has been deprecated (Anil et al., 2023) and is no longer accessible. Therefore, we use a comparable language model, gpt-4o, to conduct experiments on FetaQA and compare the results with previous methods. Additionally, we use 20 exemplars for few-shot in-context learning to align with the dataset’s format.
@@
However, the improvement of TableMaster over baseline methods remains marginal, with some values even falling below those of previous approaches in absolute terms. We believe this does not fully reflect the model’s actual performance in free-form QA. We attribute this to the n-gram text similarity metrics used in ROUGE-1/2/L (Lin, 2004), which are known to be insensitive to improvements gained from in-context learning (Maynez et al., 2023). These metrics struggle to capture stylistic and structural enhancements in free-form text generation. Since models rely on instructions and a limited number of examples, they may not fully adapt to the expected output format, leading to an underestimation of performance gains.
@@
Appendix D Table Understanding Baselines
@@
Additionally, we find that chain of thought reasoning is highly effective, achieving strong accuracy across models. Even a simple chain of thought approach outperforms previous methods that rely solely on symbolic reasoning (Mao et al., 2024), indicating that chain of thought should be retained as a key component in the reasoning framework.
@@
Appendix E Performance Analysis Under Different Table Sizes
@@
Binder (Cheng et al., 2023) 56.54 26.13 6.41
Dater (Ye et al., 2023) 62.50 42.34 34.62
Chain-of-Table (Wang et al., 2024) 68.13 52.25 44.87
@@
Table 5 presents a performance comparison across different table sizes, categorized into small (<2k tokens), medium (2$\sim$4k tokens), and large (>4k tokens). The results compare several methods, including Binder (Cheng et al., 2023), Dater (Ye et al., 2023), and Chain-of-Table (Wang et al., 2024), against TableMaster. All methods are evaluated using gpt-3.5-turbo, with additional results of TableMaster provided for gpt-4o-mini.
@@
Appendix F Performance Analysis Under Different Table Peek Sizes
@@
To analyze the effectiveness of this approach, we first examine the row count distribution in the WikiTQ dataset, as shown in Figure 6. To improve visualization, we remove 72 extreme outliers with exceptionally large row counts. The resulting density plot illustrates that the majority of tables contain fewer than 20 rows, with a pronounced peak around 10 rows. As the number of rows increases, the density gradually declines, indicating that large tables are relatively uncommon. Although a small number of tables exceed 100 rows, their frequency is minimal.
The line graph in Figure 6 illustrates the variation in accuracy with different peek sizes, where the peek size determines the number of rows considered during processing. Initially, accuracy is relatively low when only a small number of rows (e.g., 2–4) are used, reaching its minimum at a peek size of 4. We hypothesize that this occurs because, at a peek size of 2, the table includes only the top headers and a single example row, which may provide a clear structure for the language model to follow. However, at a peek size of 4, the table includes three example rows, potentially causing the language model to overfit the first few rows and misinterpret the overall table structure. This misalignment may lead to ineffective SQL generation for row lookup, resulting in a temporary drop in accuracy.
@@
Appendix G Efficiency Analysis of TableMaster
G.1 Theoretical Analysis
@@ G.2 Empirical Analysis
Recent advancements in table understanding, such as Chain-of-Table (Wang et al., 2024) and Tree-of-Table (Ji et al., 2024), involve a step-by-step evolution of tables through a long chain of transformations. In each new step, both the original table and the newly generated sub-table must be processed by language models. Additionally, their iterative process is complex, unstable, and difficult to analyze theoretically. In contrast, our approach is general and comprehensive, avoiding the trivial overhead of sub-table extraction. Instead, it focuses on holistic reasoning while maintaining ideal efficiency.
Appendix H Detailed Algorithm of Table-of-Focus Re-Construction
@@
Appendix I Analysis of Table-of-Focus Re-Construction
@@
Appendix J Analysis of Adaptive Reasoning
@@
Self-consistency methods, which perform multiple inferences to improve reliability, achieve better results. Five-shot CoT self-consistency reaches 74.98%, while five-shot PoT self-consistency lags behind at 63.97%. As introduced in (Liu et al., 2024b), mixed self-consistency (3 CoT + 3 PoT) and (5+5) further improve accuracy to 76.70% and 77.46%, respectively, at the cost of increased inference time (6 and 10 passes). Self-evaluation (self-eval) first performs CoT and PoT inferences (Prompt 26), then selects the better result, achieving 70.58% with 2 inferences.
@@
Appendix K Information Missing and Table Reasoning with Full Table
@@
PoTable (Previous SOTA) (Mao et al., 2024) 64.73 88.93
@@
To highlight the performance of TableMaster, we report the best scores of 78.13 and 90.12 in the main results table. Regardless, our method consistently outperforms the previous state-of-the-art, PoTable (Mao et al., 2024), on both WikiTQ and TabFact.
Appendix L Case Study
L.1 Case Study of Table Verbalization
@@
L.2 Case Study of TableMaster
@@
Appendix M Prompt Design in TableMaster
@@
Appendix N Prompts Used in Analysis Experiments
@@
Appendix O Notion Table

### v3

@@ 5.2 Main Results
As shown in Table 1, our TableMaster approach consistently achieves the highest performance across both WikiTQ and TabFact under different backbone models (gpt-3.5-turbo, gpt-4o-mini, and Llama-3.1-70B). On WikiTQ, TableMaster outperforms the strongest baselines by +3.51, +13.40, and +12.39 points, respectively. A similar trend is observed on TabFact, with improvements of +1.73, +1.19, and +4.10 points, demonstrating the robustness of our method across diverse large language models. Results on the FetaQA, HiTab, FinQA dataset are provided in Appendix E.1, E.2, E.3. These results confirm that TableMaster not only generalizes well across different base language models but also significantly enhances table understanding and reasoning in complex QA tasks.
@@ 5.3 Ablation Study
These results reveal that TableMaster ’s performance relies on a combination of structural understanding, content adaptation, and an effective reasoning strategy. Among these, structural extraction and reasoning components contribute the most to performance, with textual reasoning being the most critical. Additionally, while content understanding components such as re-construction and verbalization provide meaningful improvements, their impact varies across datasets. These findings validate the effectiveness of our framework in handling diverse table understanding tasks. More analysis of adaptive reasoning can be found at Appendix L.
@@ 6 Conclusion
In this paper, we explore table understanding with language models. Given the characteristics of tabular data, we identify key challenges in table understanding. To overcome these challenges, we propose TableMaster, a recipe and comprehensive framework that integrates multiple solutions. Extensive analyses and experiments demonstrate our findings and the effectiveness of TableMaster. In the future, we plan to extend and refine the framework to improve its performance across diverse practical applications, where discussed in Appendix C.
Appendix A Ethics and Broader Impacts
This work focuses on improving the table understanding capabilities of large language models through modular reasoning strategies. It does not involve the use of private, sensitive, or human-subject data. All datasets used are publicly available and widely adopted in prior work. The proposed method, TableMaster, can benefit a wide range of practical applications such as web search, spreadsheet automation, and enterprise data querying. We acknowledge that improved table reasoning systems could be misused in generating misleading summaries or automating decision-making without human oversight. However, we expect the modular and transparent nature of our method to encourage responsible deployment. No foreseeable negative societal impacts arise directly from this research.
Appendix B Scope, Key Contributions, and Comparison with Prior Work
Below we outline how TableMaster advances table understanding and how it fundamentally departs from earlier work.
A Recipe and a unifying framework, not a single algorithm.
TableMaster is proposed as a Recipe for general table understanding with language models. While most previous studies focus on a single aspect—often with specialized, heavyweight components—TableMaster supplies four lightweight yet complementary modules that together span the entire processing pipeline. This design makes TableMaster easy to adapt across diverse downstream tasks.
Challenge–solution mapping.
Section 3 and much of the appendix distill four intrinsic properties of tabular data—Structured, Intensive, Concise, Numerical—into four corresponding challenges:
  1. 1.
    Target-data localization
  2. 2.
    Semantic deficiency
  3. 3.
    Numerical inaccuracy in textual reasoning
  4. 4.
    Semantic rigidity in symbolic reasoning
For each challenge we propose a dedicated, minimal solution, whereas earlier efforts typically address only one point and overlook the others.
Efficient subtable extraction and symbolic reasoning.
Prior systems rely on elaborate heuristics for subtable extraction, which often lose information. TableMaster instead couples LLM-based column selection with SQL-style row filtering, then reconstructs a compact Table-of-Focus that preserves context—a simpler yet more general strategy.
Table verbalization for semantic enrichment.
To mitigate semantic deficiency, TableMaster inserts a Table-to-Text verbalization step that enriches input when header cues are sparse. This re-purposes a classic task to bolster table comprehension and, to our knowledge, has not been explored in prior work.
Adaptive textual–symbolic reasoning.
Although chain-of-thought reasoning has improved substantially, most earlier approaches still rely mainly on symbolic logic. We analyze the strengths and weaknesses of both paradigms (Appendix L) and introduce an adaptive scheme that dynamically chooses between—or blends—textual and symbolic reasoning without costly sampling or majority voting.
Impact.
TableMaster achieves consistent gains across five benchmarks, offers an extensible baseline for real-world deployment, and paves the way for next-generation language models equipped for robust, general-purpose table understanding.
Appendix C Limitations, Extendability, and Future Works
@@
C.1 Technical Refinement
Wild Table. In our experiments, the tables in the three datasets we use are already cleaned; therefore, we do not explicitly implement table normalization in our evaluation experiments. However, we conduct analysis experiments to highlight the importance of table normalization for handling wild tables. In practical scenarios, various tools are available for table normalization. Regular expression matching can be employed for formatting, and small language models can also be leveraged to efficiently process and normalize tables [63].
@@ C.1 Technical Refinement
Table-of-Focus Construction. In designing the Table-of-Focus, we employ two efficient methods: LM prompting for column lookup and SQL generation for row lookup. The Table-of-Focus is then constructed based on the results of these two lookups. Many previous works [45, 17] have introduced complex approaches for extracting relevant sub-tables. In contrast, our method remains intentionally simple, prioritizing efficiency and adaptability. We believe that in the future, more advanced techniques may emerge to further enhance the extraction of key information.
Table Verbalization. To facilitate the implementation of TableMaster, we utilize language models themselves to verbalize the table. However, the quality of the generated text is not optimal due to the challenges of open-ended text generation. Several existing studies, such as Table-to-Text [51], have explored this sub-task. In the future, we can enhance performance and efficiency by replacing this step with specifically trained small language models, which could further improve the semantic density of the verbalized table.
Adaptive Reasoning. Adaptive reasoning can be unstable, as language models may not always select the optimal strategy. We further explore this issue in Appendix L. In the future, training a dedicated machine learning model to guide LMs in selecting the most effective reasoning strategy could improve stability and performance.
Information Missing. The construction of the Table-of-Focus involves a trade-off between precision and recall. If recall is insufficient, essential information may be missing for final reasoning, while low precision can render the extracted content less useful. Although we use re-construction to mitigate information loss during the Table-of-Focus construction process, our analysis reveals that some information missing persist in row lookup. We further investigate this issue in Appendix M.
Efficiency. Efficiency is crucial in table processing and table understanding. To enhance efficiency, we incorporate the table peek technique, which reduces the context that language models need to process at certain steps. We further explore this technique in Appendix H and analyze the overall efficiency in Appendix I. In real-world applications, for optimal efficiency, we consider replacing certain steps with specialized small language models, balancing the trade-off between efficiency and performance .
C.2 Downstream Applications
@@
Appendix D Detailed Settings of Challenge Analysis Experiments
We conduct extensive experiments to analyze the challenges of table understanding with language models (LMs). Specifically, we perform challenge analysis experiments on the WikiTQ dataset [11], which consists of 4,344 data instances. Following previous work [17, 46], we use the exact match of the final answer as the evaluation metric to measure accuracy. Our experiments utilize OpenAI models hosted on Microsoft Azure. Unless otherwise stated, we set the temperature to 0 to ensure stable output while keeping all other hyperparameters at their default values. For each model, we use the following versions: gpt-4o (gpt-4o-0806), gpt-4o-mini (gpt-4o-mini-0718), gpt-3.5-turbo (gpt-3.5-turbo-0125), and o1 (o1-preview-0912).
@@ Appendix D Detailed Settings of Challenge Analysis Experiments
Appendix E Extended Experiments on Additional Table-Understanding Benchmarks
We perform additional experiments on diverse table-understanding tasks to further assess the robustness of TableMaster.
E.1 Evaluation on Free-form QA with the FetaQA Dataset
@@
Fine-Tuning (T5-large) [16] 30.54 63 41 53
End-to-End QA (Codex) [64] 27.96 62 40 52
End-to-End QA (PaLM 2) [17] 28.37 63 41 53
Dater (PaLM 2) [16] 29.47 63 41 53
Chain-of-Table (PaLM 2) [17] 32.61 (+14.9%) 66 (+4.8%) 44 (+7.3%) 56 (+5.7%)
@@
PaLM 2 has been deprecated [65] and is no longer accessible. Therefore, we use a comparable language model, gpt-4o, to conduct experiments on FetaQA and compare the results with previous methods. Additionally, we use 20 exemplars for few-shot in-context learning to align with the dataset’s format.
@@
However, the improvement of TableMaster over baseline methods remains marginal, with some values even falling below those of previous approaches in absolute terms. We believe this does not fully reflect the model’s actual performance in free-form QA. We attribute this to the n-gram text similarity metrics used in ROUGE-1/2/L [59], which are known to be insensitive to improvements gained from in-context learning [66]. These metrics struggle to capture stylistic and structural enhancements in free-form text generation. Since models rely on instructions and a limited number of examples, they may not fully adapt to the expected output format, leading to an underestimation of performance gains.
@@
E.2 Evaluation on Hierarchical Tables with the HiTab Dataset
UUnlike WikiTQ and TabFact, the HiTab dataset [60] contains hierarchical tables that violate the flat row–column assumption. Such structure challenges models to reason across multi-level headers and parent–child cell relations.
We employ a GPT-4o backbone and follow HiTab’s official data split. MultiCoT [67] extends Chain-of-Table reasoning to multiple tables; both MultiCoT and TableMaster operate on the same converted relational tables. E5 [68] represents the current SOTA on HiTab, being explicitly tailored for complex tables. Because our framework (TableMaster) is designed for relational tables, we introduce a lightweight relational-table converter via prompting o1: each hierarchical table is decomposed into several relational subtables, while contextual tags are propagated to preserve structural cues.
Table 4: Accuracy (%) on the HiTab dataset. “After Converting” rows evaluate models on relational tables produced by our converter; “Direct” reports results on the original hierarchical tables.
Method Accuracy
After Converting to Relational Tables
   MultiCoT (original [67]) 64.0
   MultiCoT (optimized prompt) 70.0
   MultiCoT (optimized prompt + verbalized table) 73.5
   TableMaster 74.2
Direct
   E5 [68] 77.3
Table 4 shows that TableMaster outperforms the strongest chain-of-table baseline by +0.7 point (from 73.5 to 74.2). Residual errors are mainly due to information loss during the conversion step. Future work will integrate hierarchical relations directly into the reasoning module.
E.3 Evaluation on Numerical Reasoning with the FinQA Dataset
FinQA [61] requires multi-step numerical reasoning over financial reports—e.g., computing growth rates or combining multiple cells with arithmetic operators. Hence it evaluates whether TableMaster can execute numerical formulas correctly, not just extract text spans. We keep the same training recipe as in. Two backbones are considered: GPT-4o-mini (4m) and GPT-4o (4o).
Table 5: Accuracy on FinQA. TableMaster consistently boosts numerical-reasoning accuracy over both backbones.
Method Accuracy (%) $\Delta$
GPT-4o-mini 50.7
TableMaster (4m) 66.4 +15.7
GPT-4o 63.1
TableMaster (4o) 70.9 +7.8
Table 5 shows that TableMaster delivers impressive improvements on both backbones, demonstrating that its symbolic reasoning module reliably handles complex calculations in the financial domain.
Appendix F Table Understanding Baselines
@@
Additionally, we find that chain of thought reasoning is highly effective, achieving strong accuracy across models. Even a simple chain of thought approach outperforms previous methods that rely solely on symbolic reasoning [41], indicating that chain of thought should be retained as a key component in the reasoning framework.
@@
Appendix G Performance Analysis Under Different Table Sizes
@@
Binder [18] 56.54 26.13 6.41
Dater [16] 62.50 42.34 34.62
Chain-of-Table [17] 68.13 52.25 44.87
@@
Table 7 presents a performance comparison across different table sizes, categorized into small (<2k tokens), medium (2$\sim$4k tokens), and large (>4k tokens). The results compare several methods, including Binder [18], Dater [16], and Chain-of-Table [17], against TableMaster. All methods are evaluated using gpt-3.5-turbo, with additional results of TableMaster provided for gpt-4o-mini.
@@
Appendix H Performance Analysis Under Different Table Peek Sizes
@@
To analyze the effectiveness of this approach, we first examine the row count distribution in the WikiTQ dataset, as shown in Figure 6(a). To improve visualization, we remove 72 extreme outliers with exceptionally large row counts. The resulting density plot illustrates that the majority of tables contain fewer than 20 rows, with a pronounced peak around 10 rows. As the number of rows increases, the density gradually declines, indicating that large tables are relatively uncommon. Although a small number of tables exceed 100 rows, their frequency is minimal.
The line graph in Figure 6(b) illustrates the variation in accuracy with different peek sizes, where the peek size determines the number of rows considered during processing. Initially, accuracy is relatively low when only a small number of rows (e.g., 2–4) are used, reaching its minimum at a peek size of 4. We hypothesize that this occurs because, at a peek size of 2, the table includes only the top headers and a single example row, which may provide a clear structure for the language model to follow. However, at a peek size of 4, the table includes three example rows, potentially causing the language model to overfit the first few rows and misinterpret the overall table structure. This misalignment may lead to ineffective SQL generation for row lookup, resulting in a temporary drop in accuracy.
@@
Appendix I Efficiency Analysis of TableMaster
I.1 Theoretical Analysis
@@ I.2 Empirical Analysis
Recent advancements in table understanding, such as Chain-of-Table [17] and Tree-of-Table [45], involve a step-by-step evolution of tables through a long chain of transformations. In each new step, both the original table and the newly generated sub-table must be processed by language models. Additionally, their iterative process is complex, unstable, and difficult to analyze theoretically. In contrast, our approach is general and comprehensive, avoiding the trivial overhead of sub-table extraction. Instead, it focuses on holistic reasoning while maintaining ideal efficiency.
Appendix J Detailed Algorithm of Table-of-Focus Re-Construction
@@
Appendix K Analysis of Table-of-Focus Re-Construction
@@
Appendix L Analysis of Adaptive Reasoning
@@
Self-consistency methods, which perform multiple inferences to improve reliability, achieve better results. Five-shot CoT self-consistency reaches 74.98%, while five-shot PoT self-consistency lags behind at 63.97%. As introduced in [46], mixed self-consistency (3 CoT + 3 PoT) and (5+5) further improve accuracy to 76.70% and 77.46%, respectively, at the cost of increased inference time (6 and 10 passes). Self-evaluation (self-eval) first performs CoT and PoT inferences (Prompt 26), then selects the better result, achieving 70.58% with 2 inferences.
@@
Appendix M Information Missing and Table Reasoning with Full Table
@@
PoTable (Previous SOTA) [41] 64.73 88.93
@@
To highlight the performance of TableMaster, we report the best scores of 78.13 and 90.12 in the main results table. Regardless, our method consistently outperforms the previous state-of-the-art, PoTable [41], on both WikiTQ and TabFact.
Appendix N Case Study
N.1 Case Study of Table Verbalization
@@
N.2 Case Study of TableMaster
@@
Appendix O Prompt Design in TableMaster
@@
Appendix P Prompts Used in Analysis Experiments
@@
Appendix Q Notation Table